OpenStudy (anonymous):

a water outlet can fill a 270 liter tank in one and one half hours, how long will it take to fill a bigger tank containing 1350 liters? a. 6 hrs. b 6 1/2 hrs c. 7 hrs d. 7 1/2hrs

OpenStudy (anonymous):

Hi Welcome to Openstudy! \(\sf \Large \frac{270\ L}{0.5\ hour}= \frac{1350\ L}{x\ hour}\\x\ hours= \Large \frac{1350\ L(0.5 hour)}{270\ L}\\x= ?\) evaluate (:

OpenStudy (anonymous):

okay thanks alot man

OpenStudy (anonymous):

no problem (: